# DeepMIDE: A Multivariate Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

### v1

@@ 4.2 A multivariate IDE model for multi-height space-time offshore wind forecasting
Collecting all the advection vectors across all heights yields the set $\boldsymbol{\theta}_{t}=\{\boldsymbol{\theta}_{t}^{p}\}_{p=1}^{P}$. Doing so across all time indices yields the parameter set $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{t}\}_{t=1}^{T}$. This approach, which treats the advection parameters as time- and height-varying, is in line with our local wind dynamics. However, it requires an estimation of a very large number of parameters at each time $t$ and height $p$. Instead, we propose to embed a mapping function, denoted by $\psi(\cdot)$, to represent the parameter set $\boldsymbol{\Theta}$, given streams of exogenous weather maps $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ from a regional numerical model, where $\mathbf{X}_{t}\in\mathbb{R}^{C\times\mathcal{W}}$ denotes a set of weather maps at time $t$ representing $C$ exogenous variables, with each map acquired on a bounded rectangle of $\mathbb{R}^{2}$ denoted by $\mathcal{W}$. An example of $\mathbf{X}_{t}$ in the context of our problem are the weather maps shown in Figure 2. Thus, we have $\boldsymbol{\Theta}=\psi(\mathcal{X},\boldsymbol{\Phi})$, where $\boldsymbol{\Phi}$ denote the parameters of the functional mapping, $\psi$. The redistribution kernel in (8) can now be rewritten as in (9), yielding a multivariate, nonstationary, state-dependent redistribution kernel.
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)=\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{p}t+\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{q}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (9)
@@
where $\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{p}$ denotes the estimate of the advection vector at time $t$ and height $p$.
@@
By incorporating the state-dependent kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi})$ into the formulation of (7) and approximating the double integral at $n$ spatial locations and $P$ heights as sums, the final form of our latent process model can be expressed as in (10).
$Y_{t}^{(p)}(\mathbf{s})=\sum_{\mathbf{x}=\mathbf{s}_{1}}^{\mathbf{s}_{n}}\sum_{q=1}^{P}k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)Y_{t-1}^{(q)}(\mathbf{x})+\eta_{t}^{(p)}(\mathbf{s}).$ (10)
@@ 4.3 Deep learning as a physics extractor
The remaining question is the choice of the mapping function $\psi(\cdot)$. Given the high-dimensional nature of $\mathcal{X}$, we propose a deep learning model to approximate the functional mapping $\psi(\cdot)$ and delineate the complex relationships between the exogenous image streams $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ and the advection vectors $\boldsymbol{\Theta}$. In that regard, the embedded deep learning model acts as a “physics extractor,” by learning the advection dynamics of local wind fields from the wealth of exogenous information available to the forecaster. The proposed deep learning architecture in this work fuses AlexNet (a deep CNN-based model) (Krizhevsky et al., 2012) with a custom transformer architecture to fully process the spatial and temporal information in $\mathcal{X}$. Transformer models have shown tremendous promise in computer vision applications (Vaswani et al., 2023) and hence, are well-suited to process the stream of high-dimensional weather maps.
@@
Figure 4 provides a high-level illustration DeepMIDE’s workflow, comprising two phases: an offline training stage ($t\in\{1,\ldots,\tau\}$) and an online training and forecasting stage ($t\in\{\tau+1,\ldots,T+h\}$). In the offline stage, the exogenous weather images $\{\mathbf{X}_{t}\}_{t=1}^{\tau}$ are processed through a sequence of deep CNNs to transform the high-dimensional image data into a condensed vector representation $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$. Subsequently, this sequence of vectors $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$, representing the compressed spatial features, is fed into a transformer model, which models the temporal and vertical dynamics across the spatial features, ultimately outputting the parameters $\boldsymbol{\Theta}$ that govern the advection dynamics for each corresponding time step. The CNN architecture used is shown to the right of Figure 4, whereas the transformer model architecture, along with more details about the deep learning model, are deferred to the supplementary material document appended to this manuscript.
@@
Figure 4: Illustration of the DeepMIDE framework. During both offline and online training, the inputs to the network are the exogenous image streams $\mathbf{X}_{t}$ (e.g., Pressure, Temperature, etc) (offline $t\in 1,\ldots\tau$; online $t\in{\tau+1,\ldots,T+h}$). In the offline phase, the network parameters $\boldsymbol{\Phi}$ are learned jointly with other statistical parameters $\boldsymbol{\Omega}$ by maximizing the likelihood $\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})$. For the online phase, the deep learning model will be frozen, the network parameters $\boldsymbol{\Phi}$ are fixed and the statistical parameters $\boldsymbol{\Omega}$ will be continuously updated in light of new data. The structure of the deep CNN-based AlexNet is shown to the right.
@@
We denote by $\boldsymbol{\Phi}$ the entire set of deep learning parameters (of both the deep CNNs and the transformer), whereas the remaining statistical parameters of the IDE model are denoted by $\boldsymbol{\Omega}=\{\ell_{\mathbbm{1}_{p=q}},\sigma_{\epsilon}$, $\sigma_{\eta}$, $\ell_{\epsilon}$, $\ell_{\eta}\}$. Given $\mathcal{Z}=\{\mathbf{Z}_{t}\}$ $\forall t$, $\boldsymbol{\Phi}$, and $\boldsymbol{\Omega}$, the likelihood of DeepMIDE can be compactly expressed in closed-form as in (11).
$\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=\prod_{t}\mathcal{P}(\mathbf{Z}_{t}|\boldsymbol{\Omega},\boldsymbol{\Phi}).$ (11)
@@
The exact form of $\mathcal{P}(\mathbf{Z}_{t}|\boldsymbol{\Omega},\boldsymbol{\Phi})$ is discussed in Section 4.4. For the offline stage, we jointly optimize the full parameter set $\boldsymbol{\Phi}$ and $\boldsymbol{\Omega}$ using stochastic gradient descent (Lecun et al., 1998), with backpropagation used to compute the gradient $\nabla_{\boldsymbol{\Phi},\boldsymbol{\Omega}}\mathcal{L}(\mathcal{Z}|\boldsymbol{\Phi},\boldsymbol{\Omega})$, yielding an optimal set of parameters, denoted by $\boldsymbol{\Phi}^{*}$ and $\boldsymbol{\Omega}^{*}$. For the online stage, the deep learning model will be frozen, i.e., the network parameters $\boldsymbol{\Phi}$ are kept fixed at $\boldsymbol{\Phi}^{*}$, whereas the statistical parameters $\boldsymbol{\Omega}$ will be continuously updated in light of new observations. To issue forecasts, the incoming weather maps are fed directly to the pre-trained deep learning model to predict the advection vector parameters $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{T+h}$, which are then plugged back into the multivariate IDE to output a set of multi-height space-time forecasts for $h$-steps ahead, denoted by $\hat{\mathbf{Z}}_{T+1},\ldots,\hat{\mathbf{Z}}_{T+h}$. An obvious advantage of the online/offline structure is that the deep learning model needs to be trained only once, thereby significantly reducing the computational burden in the online phase, where the model can quickly apply the pre-learnt network parameters to newly revealed weather maps.
@@ 4.4 Inference and forecasting using DeepMIDE
As in univariate IDE models, we can leverage the Kalman Filter (KF) machinery to perform inference on the process and obtain the forecasting and filtering distributions. By writing (10) in matrix format, we have:
@@
where $\mathbf{K}_{t}$ comprises all the pairwise evaluations of the redistribution kernel. From there, the forecast distribution conditioned on all past measurements is Gaussian, i.e., $\mathbf{Y}_{t}\mid\mathbf{Z}_{1:t-1}\sim\mathcal{N}\left(\mathbf{Y}_{t\mid t-1},\mathbf{P}_{t\mid t-1}\right)$ such that $\mathbf{Y}_{t\mid t-1}$ and $\mathbf{P}_{t\mid t-1}$ are expressed as in (13).
@@
$\log\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=$ $\sum_{t}\left(-\frac{mG}{2}\log 2\pi-\frac{1}{2}\log|\mathbf{P}_{t\mid t-1}|\right.$ (15)
$\left.-\frac{1}{2}\left(\mathbf{Z}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)^{\top}\left(\mathbf{P}_{t\mid t-1}\right)^{-1}\left(\mathbf{Z}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)\right),$
where ||$\cdot$|| ⋅ | is the matrix determinant. The distribution of the $h$-step latent process forecasts $\hat{\mathbf{Y}}_{T+h}\mid\mathbf{Z}_{1:T}\sim\mathcal{N}(\hat{\mathbf{Y}}_{T+h\mid T+h-1},\hat{\mathbf{P}}_{T+h\mid T+h-1})$ can be expressed as:
@@
$\hat{\mathbf{C}}_{T+h\mid T+h-1}$ $=\Sigma_{\boldsymbol{\epsilon}}+\mathbf{H}_{T+h}\hat{\mathbf{P}}_{T+h\mid T+h-1}\mathbf{H}_{T+h}^{\top}.$
@@ 5 Real-world experiments and discussions
We focus on 24-hour forecasts at $10$-min resolution. We train and evaluate DeepMIDE at the three locations where data is available: E05N, E06, and ASOW6. The data spans a total duration of approximately $8.5$ months, so we use six months for offline training, and the remainder is used for online training and forecast evaluation, carried out in a rolling fashion. For each roll, we make $24$-hour forecasts, slide by $24$ hours, and repeat the whole process. This corresponds to a number of rolls $\mathcal{R}=70$, yielding a total of $6$ forecasts/hour $\times$ $24$-hour horizon $\times$ $70$ rolls $\times$ $3$ spatial sites $\times$ $3$ heights = $90{,}720$ testing instances (In this paper, the observation mapping matrix $\mathbf{H}_{t}$ is assumed to be an identity matrix, i.e., $m=n=3,G=P=3$). The offline training takes approximetely $7$ hours using an RTX4090 GPU with $46$ GB RAM, while the online training phase takes approximately 3 minutes.
We compare DeepMIDE against four benchmarks representative of the various forecasting approaches in the wind forecasting literature and practice. All benchmark models are initially trained using the full offline data, and then updated in the online stage using newly revealed observations at each forecast roll.
  • PER: The persistence model is widely regarded as a standard forecasting benchmark, for which the performance is often highly competitive at short-term horizons. PER simply assumes that the observed wind speed at the present time will persist into the future. It is implemented separately for each height and location.
@@
  • STGP: The spatio-temporal Gaussian process model is a standard geostatistical approach which considers spatio-temporal correlations through a separable covariance function. The model is trained using the fitrgp package in MATLAB.
  • DeepAR: The Deep Autoregressive model (DeepAR) is a deep statistical method, in which a sequence of RNNs are linked in an autoregressive-based manner. We fit the DeepAR using the GluonTS package in Python, and follow the original suggestion of the authors for setting the parameters (Salinas et al., 2020).
We present the forecasting experiments in Section 5.1, followed by some in-depth discussions about the performance of DeepMIDE.

### v2

@@ 4.2 A multi-output IDE model for multi-height space-time offshore wind forecasting
Collecting all the advection vectors across $P$ heights yields the set $\boldsymbol{\theta}_{t}=\{\boldsymbol{\theta}_{t}^{p}\}_{p=1}^{P}$. Doing so across all time indices yields the parameter set $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{t}\}_{t=1}^{T}$. This approach, which treats the advection parameters as time- and height-varying, is in line with the complex dependencies observed in multi-height offshore wind speeds, as shown in Section 3.3. However, it requires an estimation of a very large number of parameters at each time $t$ and height $p$. Instead, we propose to embed a mapping function, denoted by $\psi(\cdot)$, to represent the parameter set $\boldsymbol{\Theta}$, given streams of exogenous weather maps $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ from a regional numerical model, where $\mathbf{X}_{t}\in\mathbb{R}^{C\times\mathcal{W}}$ denotes a set of weather maps at time $t$ representing $C$ exogenous variables, with each map acquired on a bounded rectangle in $\mathbb{R}^{2}$ denoted by $\mathcal{W}$. An example of $\mathbf{X}_{t}$ in the context of our problem are the weather maps shown in Figure 2. Thus, we have $\boldsymbol{\Theta}=\psi(\mathcal{X},\boldsymbol{\Phi})$, where $\boldsymbol{\Phi}$ denotes the parameters of the functional mapping, $\psi$. The redistribution kernel in (9) can now be rewritten as in (10), yielding a multivariate nonstationary kernel with functional parameters.
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)={\sigma_{k}^{2}}\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{{(p)}}t+\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{{(q)}}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (10)
@@
where $\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{{(p)}}$ denotes the estimate of the advection vector at time $t$ and height $p$.
@@
By incorporating the kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi})$ into the formulation of (8), the resulting latent process model at location $\mathbf{s}$ and height $p$ can be expressed in its final form as in (11).
$Y_{t}^{(p)}(\mathbf{s})=\sum_{\mathbf{x}=\mathbf{s}_{1}^{*}}^{\mathbf{s}_{n}^{*}}\sum_{q=1}^{P}\hat{k}^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)Y_{t-1}^{(q)}(\mathbf{x})+\eta_{t}^{(p)}(\mathbf{s}),$} (11)
where $\hat{k}^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)$ is the kernel estimated by the data-driven mapping function $\psi(\cdot)$. Note that many IDE-based models in the literature employ a reduced-rank basis expansion, which may exhibit limited performance in certain settings (STEIN2014). This issue is not as relevant in our context, since the goal is to describe the space-time dependence considering a fairly small number of observational locations. This enables us to work directly with the discretized IDE operator without relying on an explicit basis function expansion.
@@ 4.3 Deep learning as a physics extractor
The remaining question is the choice of the mapping function $\psi(\cdot)$. Given the high-dimensional nature of $\mathcal{X}$, we propose a deep learning model to approximate the functional mapping $\psi(\cdot)$ and delineate the complex relationships between the exogenous image streams $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ and the advection vectors $\boldsymbol{\Theta}$. In that regard, the embedded deep learning model acts as a “physics extractor,” by learning the advection dynamics of local wind fields from the wealth of exogenous information available to the forecaster. The proposed deep learning architecture in this work fuses AlexNet (a deep CNN-based model) (krizhevsky2012imagenet) with a custom transformer architecture to fully process the spatial and temporal information in $\mathcal{X}$. Transformer models have shown tremendous promise in computer vision applications (vaswani2023attention) and hence, are well-suited to process the stream of high-dimensional weather maps.
@@
Figure 4 provides a high-level illustration of DeepMIDE’s workflow, comprising two phases: an offline training stage ($t\in\{1,\ldots,\tau\}$) and an online training and forecasting stage ($t\in\{\tau+1,\ldots,T+h\}$), where $\tau$ is the number of timestamps in the offline training data, $T$ is the number of timestamps in the offline and online training data, and $h>0$ is the forecast horizon. In the offline stage, the exogenous weather images $\{\mathbf{X}_{t}\}_{t=1}^{\tau}$ are processed through a sequence of deep CNNs to transform the high-dimensional image data into a condensed vector representation $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$. Subsequently, this sequence of vectors $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$, representing the compressed spatial features, is fed into a transformer model, which models the temporal and vertical dynamics across the spatial features, ultimately outputting the parameters $\boldsymbol{\Theta}$ that govern the advection dynamics for each corresponding time step. The CNN architecture used is shown to the right of Figure 4, whereas the transformer model architecture, along with more details about the deep learning model, are deferred to the supplementary material document appended to this manuscript.
@@
Figure 4: Illustration of the DeepMIDE framework. During both offline and online training, the inputs to the network are the exogenous image streams $\mathbf{X}_{t}$ (e.g., Pressure, Temperature, etc) (offline $t\in 1,\ldots\tau$; online $t\in{\tau+1,\ldots,T+h}$). In the offline phase, the network parameters $\boldsymbol{\Phi}$ are learned jointly with other statistical parameters $\boldsymbol{\Omega}$ by maximizing the likelihood $\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})$. For the online phase, the deep learning model will be frozen, the network parameters $\boldsymbol{\Phi}$ are fixed and the statistical parameters $\boldsymbol{\Omega}$ will be continuously updated in light of new data. The structure of the deep CNN-based AlexNet is shown to the right. The transformer model architecture, along with more details about the deep learning model, are deferred to the supplementary material document appended to this manuscript.
@@
We denote by $\boldsymbol{\Phi}$ the entire set of deep learning parameters (of both the deep CNNs and the transformer), whereas the remaining statistical parameters of the IDE model are denoted by $\boldsymbol{\Omega}=\{\ell_{\mathbbm{1}_{p=q}},\sigma_{\epsilon}$, $\sigma_{\eta}$, $\sigma_{k}$, $\ell_{\epsilon}$, $\ell_{\eta}\}$. Given $\mathcal{Z}=\{{\mathbf{Z}}_{t}\}$ $\forall t$, $\boldsymbol{\Phi}$, and $\boldsymbol{\Omega}$, the likelihood of DeepMIDE can be compactly expressed in closed-form as in (12).
$\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=\mathcal{P}({\mathbf{Z}}_{1}|\boldsymbol{\Omega},\boldsymbol{\Phi})\prod_{t}\mathcal{P}({\mathbf{Z}}_{t}|{\mathbf{Z}}_{1:t-1},\boldsymbol{\Omega},\boldsymbol{\Phi}),$ (12)
@@
where ${\mathbf{Z}}_{1:t-1}\equiv\{{\mathbf{Z}}_{1},\ldots,{\mathbf{Z}}_{t-1}\}$. The exact form of $\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})$ is discussed in Section 4.4. For the offline stage, we jointly optimize the full parameter set $\boldsymbol{\Phi}$ and $\boldsymbol{\Omega}$ using stochastic gradient descent (sgd), with backpropagation used to compute the gradient $\nabla_{\boldsymbol{\Phi},\boldsymbol{\Omega}}\mathcal{L}(\mathcal{Z}|\boldsymbol{\Phi},\boldsymbol{\Omega})$, yielding an optimal set of parameters, denoted by $\boldsymbol{\Phi}^{*}$ and $\boldsymbol{\Omega}^{*}$. For the online stage, the deep learning model will be frozen, i.e., the network parameters $\boldsymbol{\Phi}$ are kept fixed at $\boldsymbol{\Phi}^{*}$ and no longer updated, whereas the statistical parameters $\boldsymbol{\Omega}$ will be continuously estimated to maximize the likelihood in light of new observations. To issue forecasts, the incoming weather maps are fed directly to the pre-trained deep learning model to predict the advection vector parameters $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{T+h}$, which are then plugged back into the multi-output IDE to output a set of multi-height space-time forecasts for $h$-steps ahead denoted by $\hat{\mathbf{Z}}_{T+1},\ldots,\hat{\mathbf{Z}}_{T+h}$. An obvious advantage of the online/offline structure is that the deep learning model needs to be trained only once, thereby significantly reducing the computational burden in the online phase, where the model can quickly apply the pre-learnt network parameters to newly revealed weather maps.
@@ 4.4 Inference and forecasting using DeepMIDE
As in single-output IDE models, we can leverage the Kalman Filter (KF) machinery to perform inference on the process and obtain the forecasting and filtering distributions. By writing (11) in matrix format, we have:
@@
where $\mathbf{K}_{t}$ comprises all the pairwise evaluations of the redistribution kernel in (10). The dependence of $\mathbf{K}_{t}$ on $\mathbf{X}_{t}$ is dropped for notational convenience. From there, the forecast distribution conditioned on all past measurements is Gaussian, i.e., $\mathbf{Y}_{t}\mid\ {\mathbf{Z}}_{1:t-1}\sim\mathcal{N}\left(\mathbf{Y}_{t\mid t-1},\mathbf{P}_{t\mid t-1}\right)$ such that $\mathbf{Y}_{t\mid t-1}$ and $\mathbf{P}_{t\mid t-1}$ are expressed as in (14).
@@
$\log\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=$ $\sum_{t}\left(-\frac{mG}{2}\log(2\pi)-\frac{1}{2}\log(|\mathbf{H}_{t}\mathbf{P}_{t|t-1}\mathbf{H}_{t}^{\top}+{\Sigma_{\boldsymbol{\epsilon}}}|)\right$. (16)
$\left.-\frac{1}{2}\left({\mathbf{Z}}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)^{\top}\left(\mathbf{H}_{t}\mathbf{P}_{t|t-1}\mathbf{H}_{t}^{\top}+{\Sigma_{\boldsymbol{\epsilon}}}\right)^{-1}\left({\mathbf{Z}}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)\right),$
where $|\cdot|$ is the matrix determinant. The distribution of the $h$-step latent process forecasts $\hat{\mathbf{Y}}_{T+h}\mid{\mathbf{Z}}_{1:T}\sim\mathcal{N}(\hat{\mathbf{Y}}_{T+h\mid T+h-1},\hat{\mathbf{P}}_{T+h\mid T+h-1})$ can be expressed as:
@@
$\hat{\mathbf{C}}_{T+h\mid T+h-1}$ $=\Sigma_{\boldsymbol{\epsilon}}+\mathbf{H}_{T+h}\hat{\mathbf{P}}_{T+h\mid T+h-1}\mathbf{H}_{T+h}^{\top},$
where $\hat{\mathbf{Y}}_{T+h\mid T+h-1}$ and $\hat{\mathbf{P}}_{T+h\mid T+h-1}$ denote the forecast vector of the latent state process and its covariance matrix, whereas $\hat{\mathbf{Z}}_{T+h\mid T+h-1}$ and $\hat{\mathbf{C}}_{T+h\mid T+h-1}$ denote the forecast vector of the observed measurements and its covariance matrix.
@@ 5 Real-world experiments and discussions
In this work, we focus on $2$-hour-ahead forecasts at a $10$-minute resolution. Accurate forecasts at this look-ahead horizon are used in practice to inform important wind farm and power system operations, including resource commitment, dispatch, and market participation (nyisortc). We train and evaluate DeepMIDE at the three locations where data is available: E05N, E06, and ASOW6. The data spans a total duration of approximately $8.5$ months, so we use six months for offline training, and the remainder is used for online training and forecast evaluation, carried out in a rolling fashion. Careful handling of the time index in the redistribution kernel in (10) is warranted, considering that the time index can, in theory, grow without bound leading to inconsistent physical behavior. Hence, training and forecasting are performed in rolling windows with normalized time, where $t$ is reset at the start of each window. Similarly, the advection parameters are bounded to ensure physically meaningful values. For each roll, we make $2$-hour forecasts, slide by six hours, and repeat the whole process. This corresponds to $\mathcal{R}=284$ rolls, yielding a total of $6$ forecasts/hour $\times$ 22-hour horizon $\times$ $284$ rolls $\times$ 33 spatial sites $\times$ 33 heights = 30,67230{,}672 testing instances. In this paper, the observation mapping matrix $\mathbf{H}_{t}$ is assumed to be an identity matrix, i.e., $m=n=3,G=P=3$, such that the observation and process locations are the same, that is, $\mathbf{s}_{i}=\mathbf{s}^{*}_{i}$ $\forall i=1,...,n$. This is because our goal is to obtain forecasts at the same locations where the wind turbines are installed. We find that one week (7 days) of historical data is a sufficient online training data size to balance model fitting and computational efficiency. The offline training takes approximetely $7$ hours using an RTX4090 GPU with $46$ GB RAM, while the online training phase takes approximately 3 minutes.
We compare DeepMIDE against six benchmarks that are representative of the various forecasting approaches in the wind forecasting literature and practice.
  • PER: The persistence model is widely regarded as a standard forecasting benchmark. PER assumes that the observed wind speed at the present time will persist into the future. It is implemented separately for each height and location. Despite its simplicity, the performance of PER is often highly competitive at short-term horizons.
@@
  • STGP: The spatio-temporal Gaussian process model is a prevalent geostatistical approach which considers correlations over space and time, but not height. For its covariance function, we use a convex combination of Schlater’s Lagrangian kernel (SchlatherCov) and a separable squared exponential kernel, as expressed in (19).
    $\begin{split}k^{\mathrm{STGP}}\bigl(\mathbf{s},\mathbf{x};t_{1},t_{2},\boldsymbol{\nu},\boldsymbol{\Sigma}\bigr)={}&\sigma_{k}^{2}\Biggl(\rho\,\frac{1}{\sqrt{\,\bigl|\mathbf{I}_{2\times 2}+2\,\boldsymbol{\Sigma}(t_{1}-t_{2})^{2}\bigr|\,}}\,\exp\Bigl\{-(\mathbf{s}-\mathbf{x}-\boldsymbol{\nu}(t_{1}-t_{2}))^{T}\,\\ &\bigl(\mathbf{I}_{2\times 2}+2\,\boldsymbol{\Sigma}(t_{1}-t_{2})^{2}\bigr)^{-1}\,(\mathbf{s}-\mathbf{x}-\boldsymbol{\nu}(t_{1}-t_{2}))\Bigr\}+\\ &(1-\rho)\,\Bigl(\exp\bigl(-\tfrac{\|\mathbf{s}-\mathbf{x}\|^{2}}{\ell_{\mathbf{s}}^{2}}\bigr)\,\exp\bigl(-\tfrac{\|t_{1}-t_{2}\|^{2}}{\ell_{{t}}^{2}}\bigr)\Bigr)\Biggr),\end{split}$ (19)
    where $\rho$ is an asymmetry parameter denoting the strength of asymmetry, $\ell_{\mathbf{s}}^{2}$ and $\ell_{t}^{2}$ are spatial and temporal length-scale parameters, respectively. The advection vectors in Schlather’s Lagrangian kernel are assumed to be random vectors following a bivariate Gaussian distribution for which the mean $\boldsymbol{\nu}$ and covariance $\boldsymbol{\Sigma}$ are estimated using the wind velocity forecasts from the NWP model. A separate STGP model is trained for each height using the fitrgp package in MATLAB. Hence, this space-time model, while capable of capturing complex space-time correlation structures, does not account for vertical dependencies and interactions.
  • DeepAR: The Deep Autoregressive model (DeepAR) is the space-time model proposed by deepar in which a sequence of RNNs is linked in an autoregressive-based manner. Specifically, the historical observations are fed into a deep recurrent neural network with Long Short-Term Memory (LSTM) cells, which directly outputs the parameters of a statistical likelihood function (e.g., Gaussian). Although it is fundamentally deep-learning-based (the main workhorse of making forecasts is an RNN), the relevance of this benchmark is that—similar to DeepMIDE—it falls under the general category of statistical deep learning methods. We train the DeepAR using the GluonTS package in Python, and follow the authors’ suggestions for setting the parameters (deepar).
  • DeepIDE-H: DeepIDE-H is an adaptation of the single-output DeepIDE model from zammit2020deep, where height is incorporated as an additional spatial coordinate. In that regard, DeepIDE-H combines the horizontal and vertical dimensions into a three-dimensional spatial domain. The redistribution kernel for DeepIDE-H is denoted as $k^{\text{IDE-H}}$ and is defined in (20) as the product of an asymmetric kernel for the horizontal spatial components and a squared exponential kernel for the vertical component. Here, $\boldsymbol{\theta}_{t}$ and $\gamma_{t}$ are estimated using a deep learning model that shares the same architecture as the one used in DeepMIDE. However, unlike DeepMIDE, this benchmark is a single-output model as it does not distinguish height as a physically meaningful vertical axis, and treats it as a spatial coordinate.
    $k^{\text{IDE-H}}\left(\mathbf{s},\mathbf{x};\kappa_{e},\kappa_{f},\boldsymbol{\theta}_{t},\gamma_{t}\right)=\sigma_{k}^{2}\exp\left(-\frac{\left\|\mathbf{s}-\mathbf{x}-\boldsymbol{\theta}_{t}\right\|^{2}}{\ell_{1}^{2}}\right)\exp\left(-\frac{\|\kappa_{e}-\kappa_{f}-\gamma_{t}\|^{2}}{\ell_{2}^{2}}\right),$ (20)
    where $\kappa_{e}$ and $\kappa_{f}$ are the actual altitude values for height $e$ and $f$, respectively, whereas $\ell_{1}^{2}$ and $\ell_{2}^{2}$ are length-scale parameters.
  • NWP: This is the offshore wind speed forecast from the physics-based numerical weather prediction (NWP) model: RU-WRF, which is introduced in Section 3.2. The forecasts are available in hourly resolution for each height and location.
The above six benchmarks can be grouped into the following categories representing different paradigms for wind forecasting: (i) Statistical methods: PER, ARIMA, and STGP; (ii) Statistical deep learning methods: DeepAR and DeepIDE-H; and (iii) Physics-based methods, comprising the base forecasts from the NWP model. We present the forecasting experiments in Section 5.1, followed by some in-depth discussions about the performance of DeepMIDE.